If you can boil 500g of water at 100 degrees Celsius to steam in 20 minutes, how much energy and what is the power of the burner and the cost of the electricity used at $.15/3.6e^6J
1
Expert's answer
2019-03-12T15:53:00-0400


Answer:


The energy required is the sum of 2 parts: (a) the energy required to raise the temperature of 0.5 kg of water from room temperature, assumed 20C, to 100C (delta T = 80K), (b) the enthalpy of vaporization to convert 0.5kg of water at 100C to steam at 100C.

Specific heat of water = 4.184kJ/kg-K

Enthalpy of vaporization of steam = 2257kJ/kg

So, the total energy required is:

Q = 0.5 x 4.184 x 80 + 0.5 x 2257 =167.24 + 1128.5 = 1295.74kj

Cost of electricity = 15/3.6e^6 x 1295.74 = 5.4 usd